1. 4(2y0 + 5y = 12
2. 12y + 6 - 11y - 1 = -8 + 10

Answer by minerva(22)

if that 0 was a )
then
4(2y) + 5y = 12
8y+5y=12
y=12/13
2. 12y + 6 - 11y - 1 = -8 + 10
y=-8+10-6+1
y=11-14
y=-3
i hope u forgive me if there is any mistake